In a normal distribution 35% of the items are under 40 and 10% are over 70. Find the mean and standard deviation.
1
Expert's answer
2021-07-13T13:36:23-0400

"P(Z<z_1)=0.35\to z_1=-0.3853."

"P(Z>z_2)=0.10\to P(Z<z_2)=0.90\to z_2=1.2816."

"\frac{40-\mu}{\sigma}=-0.3853."

"\frac{70-\mu}{\sigma}=1.2816."

"\mu-0.3853\sigma=40."

"\mu+1.2816\sigma=70."

"\sigma=\frac{70-40}{1.2816+0.3853}\approx18."

"\mu=40+0.3853*18\approx47."